OpenStudy (anonymous):

Use basic identities to simplify the expression. cos^2 theta/sin^2theta +csc theta sin theta

OpenStudy (anonymous):

( . Y . ) =]]

OpenStudy (anonymous):

what is your problem dude @sourwing

OpenStudy (anonymous):

Can you answer my question or not? @sourwing

OpenStudy (anonymous):

(sin theta) (sec theta) (cot theta) (csc theta) Can also be written as: (sin theta) 1/(cos theta) (cos theta)/(sin theta) 1/(sin theta) Many of the terms in the numerator and denominator cancel each other out, so you are left with: 1/(sin theta)=csc theta :D Whenever you have sec, csc, or cot, it helps to write them in terms of sin and cos, it makes it easier to see :D this what i can do

OpenStudy (anonymous):

There are 6 basic trig functions. sin(x) = 1/csc(x) cos(x) = 1/sec(x) tan(x) = sin(x)/cos(x) or 1/cot(x) csc(x) = 1/sin(x) sec(x) = 1/cos(x) cot(x) = cos(x)/sin(x) or 1/tan(x) In your problem csc(x)*cot(x) we can simplify csc(x). csc(x) = 1/sin(x) Similarly, cot(x) = cos(x)/sin(x). csc(x)*cot(x) = (1/sin[x])*(cos[x]/sin[x]) = cos(x)/sin2(x) = cos(x) * 1/sin2(x) Either of the above answers should work. In general, try converting your trig functions into sine and cosine to make things simpler.

OpenStudy (anonymous):

thanks for the medal

OpenStudy (anonymous):

Would it be cos y?

OpenStudy (anonymous):

@lupita1995

OpenStudy (anonymous):

give me a min

OpenStudy (anonymous):

csc2θ sec2θ 1 tan2θ These are my options